OpenStudy (anonymous):

what intervals on the graph f(x) = x^4 - 6x^2 convave upward? and which intervals concave downward? whic are inflection points?

OpenStudy (anonymous):

sorry for the mis type...it should say cancave upward..

OpenStudy (anonymous):

\[f \prime \left( x \right)=4x ^{3}-12x\] \[f \prime \left( x \right)=0 gives 4x ^{3}-12x=0,4x \left( x ^{2}-3 \right)=0\] \[x \left( x-\sqrt{3} \right)\left( x+\sqrt{3} \right)=0\] \[x=0, \sqrt{3},-\sqrt{3}\] \[f'\left( x \right)=4x \left( x-\sqrt{3} \right)\left( x+\sqrt{3} \right)\] 1. when x <0 slightly, f'(x)=-(-)(+)=+ve when x >0 f'(x)=+(-)(+)=-ve f'(x) changes sign from +ve to -ve as x passes through 0 hence there is relative maxima at x=0 \[ 2. when x<\sqrt{3}slightly\],\[x-\sqrt{3}<0\] f'(x)=+(-)(+)=-ve \[when x> \sqrt{3},x-\sqrt{3}>0\] f'(x)=+(+)(+)=+ve therefore f'(x) changes sign from -ve to +ve as x passes through \[\sqrt{3}\] Hence there is minima at x=sqrt3 \[3.when x <- \sqrt{3} slightly, x+\sqrt{3}<0\] f'(x)=-(-)(-)=-ve \[when x>-\sqrt{3},x+\sqrt{3}>0\] f'(x)=-(-)(+)=+ve Therefore f'(x) changes sign from -ve to +ve as x passes through \[-\sqrt{3}\] Hence there is local minima at x=-sqrt3 at maxima concave downward. At minima concave upward.

OpenStudy (anonymous):